Operation Summary for May 21, 2010 - May 24, 2010 |
The Integrated Luminosity for the period from 5/17/10 to 5/24/10 was 51.74 pb -1. NuMI reported receiving 7.23E18 protons on target during this same period. Friday May 21 The day shift began with Operations monitoring store 7827 (78.27E30) and the Recycler stash (221.4E10), with the Antiproton Source (Pbar) stacking (13.8E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Booster personnel worked on a RF station (BRF18), which required a short downtime. During the evening shift, Operations terminated store 7827 at 4:10 PM. Operations put the TeV into HEP shot setup and established store 7828 at 5:45 PM with an average initial luminosity (AIL) of 359.7E30. Saturday May 22 The midnight shift began with Operations monitoring store 7828 (137.02E30) and the stash (184.8E10), with Pbar stacking (23.4E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, at 4:43 AM, store 7828 quenched at TeV sector A3 due to a heater firing unit (HFU) pre-fire. During the day shift, quench recovery and TeV checkout continued until shot setup began at 9:03 AM. Operations established store 7830 at 10:08 AM with an AIL of 351.36E30. During the evening shift, Operations reported trouble with a MI-60 power supply PLC and a Booster kicker thyratron that kept beam off from 8:45 PM to 10:22 PM. Sunday May 23 The midnight shift began with Operations monitoring the store 7830, the stash, and the Pbar stack, with Booster down due to a kicker (MKS08) problem, and with NuMI and MiniBooNE waiting for the Booster beam. During the midnight shift, Operations terminated store 7830 at 4 AM. Operations put the TeV into shot setup, but at 5:13 AM, the setup ended with a TeV quench due to bad beam positions at D0. During the day shift, the quench investigation and recovery ran throughout the shift due to a Kautzky valve that had to be replaced. Operations reported at 9:54 AM that trouble with a MI-50 power supply PLC kept the MI beam off. During the evening shift, Operations reported that they brought MI online at 5:40 PM. Then the MI RF north anode power supply tripped off at 6:37 PM and kept MI beam off until 10:50 PM. And finally, Operations reported that a bad TeV A4 reciprocating cold compressor held off the TeV turn on. Monday May 24 The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations reported that the TeV turn on began at 3:10 AM. Operations put the TeV into shot setup at 5:38 AM and established store 7834 at 6:35 AM with an AIL of 333.1E30. The Plans for Monday The plans for today are to stack, stash, and store. The plans for tomorrow are to allow Pbar a six-hour access, beginning at 7 AM, to replace a dry engine flywheel and to work on the closed loop LCW system for their dump. |
|
More Information Comments and Suggestions |
| last modified 05/24/2010 email Fermilab |
| Security, Privacy, Legal | |